# CroCoDai: A Stablecoin for Cross-Chain Commerce

### v1

@@ 6.1 Smart Contracts
CCTransfer. This contract stores cross-chain transfer requests as a hashmap of Request instances, which are implemented as a Solidity struct. Each CCTransfer contract is linked to a specific Coin and Relay contract pair, and must be made a Coin contract ward to allow it to mint or burn stablecoins upon the commitment of cross-chain transfers.
@@
Relay. In the first design, a hashmap of relay node addresses in Solidity’s address format is stored in the contract. For each cross-chain transfer, at least $12f+$ 1 relay nodes must call the Relay node’s approve function using their stored address. This design exploits the EVM’s built-in signature verification, and generalizes trivially to other blockchain platforms. However, the downside is that each relay node must have an account and spend gas on each supported blockchain for every cross-chain transfer. In the second design, a hashmap of relay node public keys is stored in the contract, and a batch of $12f+$ 1 32-byte signatures is passed in the form of a byte array as a function parameter for each transaction. These signatures are then validated one by one by repeatedly calling the built-in ecrecover function. Although this is not as generic as the previous approach (e.g., each chain with its different cryptographic primitives would require its own off-chain signature generation script) it is more efficient and does not require each individual relay node to maintain an account. In the third design, we use a $t$-out-of-$n$ threshold signature scheme (where in our case $t=2f+1$) to produce a single 32-byte signature for the relay node quorum, e.g., using the procedure of [23]. The resulting signature is a standard ECDSA signature that can be verified using ecrecover. For our implementation we have used the multi-party ECDSA library by ZenGo X [59].
Table 3: Performance comparison of the three implementation designs.
@@ 6.2 Experimental Results
Table 3 compares the different designs in terms of the gas cost and the processing time at the relay nodes, $T_{\textnormal{proc}}$. We observe that the first design is the worst in terms of performance, which is the cost of its generality and ease of implementation. The second design is more efficient in terms of gas cost for large relay node quora, and only has little processing time at the relay nodes. The third design is the most efficient in terms of gas cost, but at the cost of considerable processing time at the relay nodes, especially for large quora. For the first ($n/n$) design, we could not distinguish $T_{\textnormal{proc}}$ from the transaction processing time in Ganache as both happen in a single step in Python’s Web3 package – however, these are assumed to be similar to signing in the $n/1$ design, and are indicated with ‘—’ in the table.
By comparison, zkBridge [54] reports a gas cost of $\approx$230k gas for on-chain verification, but with a processing cost of $\approx$18 seconds to generate a proof with 32 high-end Amazon Web Services machines (c5.24xlarge). This lower performance is to be expected, as the additional complexity of zkBridge removes the need for a trusted quorum of relay nodes.
7 Conclusions and Future Work
In this work we have presented CroCoDai, a stablecoin system for cross-chain commerce. CroCoDai facilitates cross-chain transfers by design, and is more resilient than a single-chain system to black swan events and price fluctuations. Our empirical analysis shows that with sufficient overcollateralization, the probability of system failure can be made small even under a model whose price data includes the highly volatile period after the collapse of the FTX exchange. We have presented three prototype implementations that rely on a quorum of relay nodes for cross-chain transfers and governance decisions.
For future work, a first interesting direction is to implement a version of CroCoDai that performs all relay operations on-chain, as per the second design of Section 4.4. Promising directions to achieve such a design would be to use trustless bridges as proposed in zkBridge, or succinct proofs of state such as those proposed for Algorand or Mina. This would remove the need for a trusted quorum of relay nodes, while maintaining CroCoDai’s advantages over the single-chain DSS, i.e., higher resilience to black swan events and price fluctuations. A second interesting direction is to further investigate alternative collateral liquidation mechanisms – for example, Salehi et al. [48] propose to mitigate liquidations from the DSS or remove them completely. For example, collateral of the same token type in a vault could be pooled together and liquidated only when the entire pool is underwater – however, this would remove an incentive from users to overcollateralize. A third and final interesting research direction is system recovery after a 51% attack on a coin chain. As we discussed in Section 5, such an attack could create stablecoins without matching collateral on a different chain. This is similar to the situation after a failed collateral auction in the DSS, so it may be possible to resolve it similarly – i.e., with the governance token holders absorbing the losses.
Appendix A List of Symbols
A list of symbols used in this paper can be found in Table 4.
Table 4: List of symbols.
symbol description
$p_{m\,t}$ base price of token $m$ at time $t$
$p_{m\,o\,t}$ price of token $m$ at time $t$, as reported by oracle $o$
$\mu_{m}$ mean drift of the stochastic process $p_{m\,t}$
$\sigma^{2}_{m}$ variance of the stochastic process $p_{m\,t}$
$\mu_{m\,o}$ mean of the random variable $p_{m\,o\,t}-p_{m\,t}$
$\sigma^{2}_{m\,o}$ variance of the random variable $p_{m\,o\,t}-p_{m\,t}$
$y_{j}$ collateral token type of CDP $j$
$c_{j\,t}$ number of collateral tokens held in CDP $j$ at time $t$
$D_{j\,t}$ debt of CDP $j$ at time $t$
$\zeta_{m}$ debt ceiling of token type $m$
$\zeta^{\prime}$ total debt ceiling of the failed/compromised asset types
$\gamma$ liquidation ratio
$\lambda$ maximum contribution of each individual asset type during convex optimization
$\theta$ system failure threshold
$\mathcal{J}_{t}$ set of all CDPs at time $t$
$M$ number of tokens
$O$ number of price feeds
$U$ number of users
Appendix B Alternative Simulation Methods
In the following, we discuss two alternative methods for estimating failure risks through simulation: using pure historical data (without an underlying price model), and with a standard GBM as discussed in Section 3.2.
For the first method, we assume for each time slot that our portfolio had a given overcollateralization at the start of the slot and check if system failure occurs within a period of interest (e.g., one day). We repeat this process for all time steps in our dataset. The downside of this approach is that our dataset captures a period with many sharp price drops, which results in a somewhat pessimistic outlook on the failure risk – on the other hand, it captures realistically the limitations of stablecoins that rely on crypto-assets.
For the second method, we generate correlated, normally distributed samples that have the same covariance matrix as the original prices. To achieve this, let $\mathbf{L}$ be computed as $\textnormal{Cov}(\mathbf{X})=\mathbf{L}\mathbf{L}^{T}$ using Cholesky decomposition. Random samples can then be generated as $\mathbf{LW}=\mathbf{R}$, where the matrix $\mathbf{W}$ contains random values sampled from the standard normal distribution. We can then calculate the returns of the assets in portfolio as
$S_{t}=S_{0}\cdot\left\{\left(\mu-\frac{\sigma^{2}}{2}\right)+\sigma W_{t}\right\}.$ (5)
We emphasize that the method used in Section 5.2 is the same as the second method, with the exception being that we use the multivariate $t$-distribution to represent the returns instead of the multivariate normal distribution.
@@ Appendix B Alternative Simulation Methods
Table 5: Estimated portfolio failure probabilities - historical method.
single-asset multi-asset $\gamma^{\prime}$     BTC     ETH     ADA     SOL     DOT     TRX     AVAX     ALGO     EOS     D-All     D-ERC     D-Mix1     D-Mix2     C-Mix1     C-Mix2 1.2 0.052 $\pm$0.0057 0.072 $\pm$0.0067 0.060 $\pm$0.0062 0.198 $\pm$0.0103 0.048 $\pm$0.0056 0.040 $\pm$0.0051 0.134 $\pm$0.0088 0.135 $\pm$0.0089 0.089 $\pm$0.0074 $0$$\pm$0.0000 0.067 $\pm$0.0065 0.060 $\pm$0.0062 0.060 $\pm$0.0062 0.058 $\pm$0.0061 0.052 $\pm$0.0058 1.3 0.002 $\pm$0.0012 0.024 $\pm$0.0040 0.003 $\pm$0.0013 0.054 $\pm$0.0059 0.006 $\pm$0.0020 0.002 $\pm$0.0012 0.020 $\pm$0.0036 0.037 $\pm$0.0049 0.022 $\pm$0.0038 $0$$\pm$0.0000 0.024 $\pm$0.0039 0.022 $\pm$0.0038 0.024 $\pm$0.0039 0.019 $\pm$0.0035 0.006 $\pm$0.0020 1.4 $0$$\pm$0.0000 0.004 $\pm$0.0016 $0$$\pm$0.0000 0.037 $\pm$0.0049 $0$$\pm$0.0000 $0$$\pm$0.0000 0.003 $\pm$0.0015 0.019 $\pm$0.0035 $0$$\pm$0.0000 $0$$\pm$0.0000 0.001 $\pm$0.0010 0.001 $\pm$0.0010 0.001 $\pm$0.0010 $0$$\pm$0.0000 $0$$\pm$0.0000 1.5 $0$$\pm$0.0000 $0$$\pm$0.0000 $0$$\pm$0.0000 0.030 $\pm$0.0044 $0$$\pm$0.0000 $0$$\pm$0.0000 $0$$\pm$0.0000 0.003 $\pm$0.0015 $0$$\pm$0.0000 $0$$\pm$0.0000 $0$$\pm$0.0000 $0$$\pm$0.0000 $0$$\pm$0.0000 $0$$\pm$0.0000 $0$$\pm$0.0000
Table 6: Estimated portfolio failure probabilities - standard GBM.
single-asset multi-asset $\gamma^{\prime}$     BTC     ETH     ADA     SOL     DOT     TRX     AVAX     ALGO     EOS     D-All     D-ERC     D-Mix1     D-Mix2     C-Mix1     C-Mix2 1.2 0.024 $\pm$0.0005 0.046 $\pm$0.0007 0.043 $\pm$0.0006 0.147 $\pm$0.0011 0.036 $\pm$0.0006 0.024 $\pm$0.0005 0.073 $\pm$0.0008 0.124 $\pm$0.0010 0.069 $\pm$0.0008 $0$$\pm$0.0000 0.045 $\pm$0.0007 0.037 $\pm$0.0006 0.042 $\pm$0.0006 0.034 $\pm$0.0006 0.031 $\pm$0.0005 1.3 0.001 $\pm$0.0001 0.005 $\pm$0.0002 0.003 $\pm$0.0002 0.070 $\pm$0.0008 0.002 $\pm$0.0001 0.001 $\pm$0.0001 0.010 $\pm$0.0003 0.025 $\pm$0.0005 0.007 $\pm$0.0003 $0$$\pm$0.0000 0.004 $\pm$0.0002 0.002 $\pm$0.0001 0.003 $\pm$0.0002 0.002 $\pm$0.0001 0.001 $\pm$0.0001 1.4 $0$$\pm$0.0000 $0$$\pm$0.0000 $0$$\pm$0.0000 0.038 $\pm$0.0006 $0$$\pm$0.0000 $0$$\pm$0.0000 0.001 $\pm$0.0001 0.004 $\pm$0.0002 $0$$\pm$0.0001 $0$$\pm$0.0000 $0$$\pm$0.0000 $0$$\pm$0.0000 $0$$\pm$0.0000 $0$$\pm$0.0000 $0$$\pm$0.0000 1.5 $0$$\pm$0.0000 $0$$\pm$0.0000 $0$$\pm$0.0000 0.019 $\pm$0.0004 $0$$\pm$0.0000 $0$$\pm$0.0000 $0$$\pm$0.0000 $0$$\pm$0.0001 $0$$\pm$0.0000 $0$$\pm$0.0000 $0$$\pm$0.0000 $0$$\pm$0.0000 $0$$\pm$0.0000 $0$$\pm$0.0000 $0$$\pm$0.0000
Table 7: Estimated portfolio failure probabilities - Student’s $t$ GBM.
single-asset multi-asset $\gamma^{\prime}$     BTC     ETH     ADA     SOL     DOT     TRX     AVAX     ALGO     EOS     D-All     D-ERC     D-Mix1     D-Mix2     C-Mix1     C-Mix2 1.2 0.022 $\pm$0.0005 0.047 $\pm$0.0007 0.077 $\pm$0.0008 0.227 $\pm$0.0013 0.062 $\pm$0.0008 0.019 $\pm$0.0004 0.126 $\pm$0.0010 0.164 $\pm$0.0012 0.099 $\pm$0.0009 0.001 $\pm$0.0001 0.050 $\pm$0.0007 0.040 $\pm$0.0006 0.048 $\pm$0.0007 0.028 $\pm$0.0005 0.029 $\pm$0.0005 1.3 0.003 $\pm$0.0002 0.008 $\pm$0.0003 0.014 $\pm$0.0004 0.092 $\pm$0.0009 0.012 $\pm$0.0003 0.003 $\pm$0.0002 0.029 $\pm$0.0005 0.051 $\pm$0.0007 0.015 $\pm$0.0004 $0$$\pm$0.0000 0.007 $\pm$0.0003 0.005 $\pm$0.0002 0.007 $\pm$0.0003 0.003 $\pm$0.0002 0.003 $\pm$0.0002 1.4 0.001 $\pm$0.0001 0.003 $\pm$0.0002 0.005 $\pm$0.0002 0.059 $\pm$0.0007 0.004 $\pm$0.0002 0.001 $\pm$0.0001 0.009 $\pm$0.0003 0.019 $\pm$0.0004 0.005 $\pm$0.0002 $0$$\pm$0.0000 0.002 $\pm$0.0001 0.001 $\pm$0.0001 0.002 $\pm$0.0001 0.001 $\pm$0.0001 0.001 $\pm$0.0001 1.5 0.001 $\pm$0.0001 0.002 $\pm$0.0001 0.002 $\pm$0.0001 0.042 $\pm$0.0006 0.002 $\pm$0.0001 0.001 $\pm$0.0001 0.004 $\pm$0.0002 0.008 $\pm$0.0003 0.002 $\pm$0.0002 $0$$\pm$0.0000 0.001 $\pm$0.0001 0.001 $\pm$0.0001 0.001 $\pm$0.0001 0.001 $\pm$0.0001 0.001 $\pm$0.0001
We visualize the difference between the standard and $t$-distribution GBM for 10 separate simulation runs in Figure 4(a) and Figure 4(b), respectively. The $t$-distribution GBM runs are visibly more likely to experience sharp jumps than the standard GBM runs.
Figure 4: For 10 separate asset price simulation runs, the total value of the portfolio 50% BTC and 50% ETH) over a 1 day period, using the standard GBM (a) and the GBM with the $t$-distribution (b). The dashed line represents the system failure threshold for $\gamma$ = 1.2 and $\theta$ = 1.1.
Appendix C Distribution Fitting Examples
The GBM is usually assumed to follow a normal distribution. In previous simulations we used normal distributions to simulate the returns. As mentioned by many papers, distribution of stock returns is usually more heavy-tailed and Student’s t-distribution is better fit (see, e.g., [34, 29]). When comparing residual sum of squares for normal and t-distribution on returns of ETH, we get a residual sum of squares (RSS) that is approximately equal to 31900 and 14000 for the normal and $t$- distribution, respectively. Moreover, the fitted $t$-distribution has a lower RSS for all other assets as well. In Figure 5, we have displayed a comparison of the normal and $t$-distribution fitted on Ethereum and Solona returns. As such, we prefer to use the $t$-distribution for the risk analysis in this paper.
Figure 5: Comparison between t-distribution and normal distribution fitted on historical returns.
Appendix D Proofs of Theorems 5.1 – 5.3
@@ Appendix C Distribution Fitting Examples
Before we prove the main theorems, we need the following technical lemmas. We need the first to prove that if all CDPs are over-collateralized, then so is system as a whole.
Lemma D.1.
@@
Proof D.2.
Note that $\frac{\sum_{i=1}^{n}a_{i}}{\sum_{i=1}^{n}b_{i}}=\sum_{i=1}^{n}\frac{a_{i}}{b_{i}}\frac{b_{i}}{\sum_{j=1}^{n}b_{j}}$. Since $\sum_{i=1}^{n}\frac{b_{i}}{\sum_{j=1}^{n}b_{j}}=1$, $\frac{\sum_{i=1}^{n}a_{i}}{\sum_{i=1}^{n}b_{i}}$ is a convex combination of $a_{i}/b_{i},\ldots,a_{n}/b_{n}$. As such, it cannot be smaller than $\min_{i=1,\ldots,n}a_{i}/b_{i}$, which is greater than $c$ by the statement of the lemma.
@@
Proof D.5 (Proof of Theorem 5.1).
If $p_{y_{j}\,t}c_{j\,t}/y_{j\,t}>\gamma$ holds for all $j\in\mathcal{J}$, then by Lemma D.1 it must hold that
$\frac{C^{+}_{0}}{D^{+}_{0}}>\gamma.$ (6)
@@
$D^{-}_{0}<\zeta^{\prime}/(1-\zeta^{\prime})D^{+}_{0}.$ (7)
@@
where the first inequality holds because of (7) and the second because of (LABEL:eq:unaffected_over-collateralized).
Proof D.6 (Proof of Theorem 5.2).
@@
$D^{-}_{0}+h^{\prime}<\zeta^{\prime}/(1-\zeta^{\prime})D^{+}_{0}$ (8)
@@
where the first inequality holds because of (8) and the second because $C^{-}_{0}>0$ and because of (6). As such, we again find that the debt ceilings ensure that CroCoDai remains fully backed if it is sufficiently over-collateralized.
Proof D.7 (Proof of Theorem 5.3).
In the following, we first assume that $O$ is odd and that some subset $O^{\prime}\subset[O]$ is controlled by malicious parties who aim to convince a vault in CroCoDai that the price is at most than some $p^{*}>p$. In this case, the median equals the $\frac{1}{2}(M+1)+|O^{\prime}|$ th order statistic from a sample of $M-|O|$ normally distributed random variables with mean $p$ and standard deviation $\sigma_{o}^{2}$. This order statistic is greater than the maximum of the $M-O^{\prime}$ random variables, which we denote by $Y$. Let $\sigma_{*}=\min_{o\in[O]}\sigma_{o}$. $Y$’s cumulative distribution function equals
$\begin{split}\mathbb{P}(Y\geq p^{*})&=\mathbb{P}\left(\bigwedge_{o\in[O]\backslash M^{\prime}}\Delta_{o}>p^{*}-p\right)=\prod_{o\notin O^{\prime}}R\left(\frac{p^{*}-p}{\sigma_{o}}\right)\\ &\leq R^{M-|O|}\left(\frac{p^{*}-p}{\sigma_{*}}\right)\leq\frac{e^{-(M-|O^{\prime}|)(p^{*}-p)^{2}/2}}{(p^{*}-p)\sigma\sqrt{2\pi}},\end{split}$
where the first inequality holds because $R$ is a monotonically decreasing function, and the second inequality follows from Lemma D.3. As such, the probability $\mathbb{P}(Y\geq p^{*})$ vanishes exponentially as $p^{*}-p$ increases. The proof for $p^{*}<p$ is very similar and uses the symmetry of the normal distribution.
Appendix E DSS Parametrization
The parameters of the various different vault types in the DSS can be found here. A sample of parameter values as of 25 April 2023 (Ethereum block 17121982) can be found in Table 8. For each of the supported token type $j$, we indicate the token name, total value of the collateral of CDPs of this type $\sum_{j\in\mathcal{J}}p_{j\,t}c_{j\,t}$, total debt of the collateral of CDPs of this type $\sum_{j\in\mathcal{J}}D_{j\,t}$, collateralization ratio, and debt ceiling $\zeta_{j}$. Although the exact values vary throughout the day, this table is meant to illustrate the typical composition of the collateral type in the DSS.

### v4

@@ 6.1. Smart Contracts
CCTransfer. This contract stores cross-chain transfer requests as a hashmap of Request instances, which are implemented using a Solidity struct. Each CCTransfer contract is linked to a specific Coin and Relay contract pair, and must be made a Coin contract ward to allow it to mint or burn stablecoins upon the commitment of cross-chain transfers.
@@
Relay. In the first design (named $n$/$n$), a hashmap of relay node addresses in Solidity’s address format is stored in the contract. For each cross-chain transfer, at least $12f+$ 1 relay nodes must call the Relay contract’s approve function using their stored address. This design exploits the EVM’s built-in signature verification, and generalizes trivially to other blockchain platforms. However, the downside is that each relay node must have an account and spend gas on each supported blockchain for every cross-chain transfer. In the second design ($n$/$1$), a hashmap of relay node public keys is stored in the contract, and a batch of $12f+$ 1 32-byte signatures is passed in the form of a byte array as a function parameter for each transaction. These signatures are then validated one by one by repeatedly calling the built-in ecrecover function. Although this is not as generic as the previous approach (e.g., each chain with its different cryptographic primitives would require its own off-chain signature generation script) it is more efficient and does not require each individual relay node to maintain an account. In the third design ($1$/$1$), we use a $t$-out-of-$n$ threshold signature scheme (where in our case $t=2f+1$) to produce a single 32-byte signature for the relay node quorum, e.g., using the procedure of (gennaro2018fast, ). The resulting signature is a standard ECDSA signature that can be verified using ecrecover. For our implementation we have used the multi-party ECDSA library by ZenGo X (zengox, ).
Table 3. Performance comparison of the three implementation subvariants.
@@ 6.2. Relay Microbenchmarks
Table 3 compares the different designs in terms of the gas cost and the processing time at the relay nodes, $T_{\textnormal{proc}}$. Both are essential to system performance: if the on-chain cost is too high then each transfer will incur large fees, whereas if the computation time is too high then cross-chain transactions incur a large delay. For the experiments, we used a Lenovo IdeaPad with an Intel(R) Core(TM) i5-1035G1 CPU (1.00GHz) and 8GB RAM – however, the gas costs are machine-independent. The scripts that performed off-chain computations use Python’s Web3 package for cryptographic operations, and the underlying blockchain was simulated using Truffle’s Ganache tool (ganache, ). To execute our experiments, we ran the Python scripts in different terminals that each interacted with an active Ganache instance on a single machine. From Table 3, we observe that the first subvariant is the worst in terms of performance, which is the cost of its generality and ease of implementation. The second subvariant is more efficient in terms of gas costs for large quora, and only has little processing time at the relay nodes. The third subvariant is the most efficient in terms of gas costs, but at the cost of considerable processing time at the relay nodes, especially for large quora. For the first subvariant (i.e., $n/n$), we were unable to distinguish $T_{\textnormal{proc}}$ from the transaction processing time in Ganache as both happen in a single step in Python’s Web3 package – however, these are assumed to be similar to signing in the $n/1$ design, and are indicated by ‘—’ in the table.
6.3. End-to-End Experiments
As discussed in Section 4.6, a baseline that uses the existing DSS with a greater variety of wrapped tokens has similar resilience to black swan events and price fluctuations as CroCoDai. However, we provide experimental results in this section that demonstrate that the DSS baseline has worse end-to-end performance than CroCoDai. We focus on the two main stablecoin actions: (1) creating new stablecoins on a coin chain and (2) transferring stablecoins from one coin chain (Chain A) to another (Chain B). To ensure that gas costs are comparable, we implement Chain A, Chain B, and the relay chain as private Ethereum chains using geth. For collateral, we implement a custom ERC20 token contract on the underlying chain (GemCoin) – for the DSS baseline, we also require a bridge for the custom token on Chain A (GemBridge), and another (DaiBridge) for Dai on the relay chain. For the collateral vault, we use instances of Vat, DaiJoin, and GemJoin where the latter is linked to the GemCoin contract – in CroCoDai the vault contracts exist on Chain A, whereas in the DSS baseline they exist on the relay chain. Finally, the baseline requires a WrGemCoin contract on the relay chain to implement the wrapped collateral token as an ERC20 token, and a WrDai contract for wrapped Dai tokens on each coin chain.
Creating Stablecoins. For CroCoDai, the procedure for creating new stablecoins is as follows: the user first transfers the collateral to the vault and creates a CDP using GemCoin and GemJoin. Next, she uses the CDP to create stablecoins and transfers them to her account using Vat and DaiJoin.
For the DSS baseline, the procedure is as follows. The user first sends GemCoin tokens to the bridge using GemCoin and BaseBridge. Next, a relay detects the transfer and creates an equivalent amount of wrapped GemCoin tokens for the user using WrGemCoin. The user then transfers the collateral to the vault, creates a CDP, creates stablecoins, and transfers them to her account using WrGemCoin, GemJoin, Vat and DaiJoin. Next, the user sends the Dai tokens to the bridge using Dai and DaiBridge. Finally, a relay detects the transfer and mints an equivalent amount of wrapped Dai tokens for the user using WrDai.
Cross-Chain Transfers. For CroCoDai, the procedure for transferring stablecoins from Chain A to B is as follows. The user first transfers the tokens to CCTransfer and creates a transfer request. Next, a relay detects the request, sends a signed approval to the relay chain and Relay on Chain A, which calls Coin to burn the stablecoins. Finally, the relay sends the signed message to Relay on Chain B, which calls Coin to mint the stablecoins.
For the DSS baseline, the procedure is as follows: the user first burns her wrapped Dai tokens using WrDai. Upon detection by a relay, the burned Dai is refunded to the user on the relay chain. The user then transfers the Dai to the DaiBridge for Chain B. Finally, a relay detects the transfer and mints an equivalent amount of wrapped Dai tokens for the user using WrDai on Chain B.
Table 4. Computation costs (units of gas) of the transactions in CroCoDai and the DSS baseline, grouped by block. Bold: relay chain computations.
CroCoDai DSS baseline
creation transfer creation transfer
Block 1 (Chain A) 466537 229325 244036 27762
Block 2 (Relay) 133280 770724 294016
Block 3 (Chain B) 185503 71498
Block 3’ (Chain A) 71498
Total (coin chains) 466537 414828 315534 99260
Total (relay chain) 0 133280 770724 294016
Results. Table 4 displays the cost of the processes discussed above, with transactions that can occur within the same block grouped together. For the DSS baseline, we assume that a single trusted custodian acts as a relay – to keep the comparison with CroCoDai as fair as possible, we hence use the first prototype subvariant ($n/n$) but with a single signer. For our implementation, we assume that the recipient data for outgoing cross-chain transfers (i.e., the ID of the receiving chain and the recipient’s account data) is not written to the blockchain; even in the fully on-chain design (see Section 3.3), this data can be exchanged off-chain and incorporated in the commit message or zero-knowledge proof.
@@ 6.3. End-to-End Experiments
7. Conclusions and Future Work
We have presented CroCoDai, a stablecoin for cross-chain commerce. CroCoDai facilitates cross-chain transfers by design, and is more resilient than a siloed baseline to black swan events and price fluctuations. We have presented a prototype implementation that relies on a quorum of relay nodes for cross-chain transfers and governance decisions, and which has better performance than a DSS-based baseline in terms of (expensive) transactions on the relay chain. For future work, an interesting direction is to implement a version of CroCoDai that performs all relay operations on-chain, as per the second variant of Section 3.3 – e.g., using the approach of zkBridge, or succinct proofs of state as in Algorand or Mina. Another interesting direction is to further investigate alternative collateral liquidation mechanisms – for example, Salehi et al. (salehi2021red, ) propose to mitigate liquidations from the DSS or remove them completely. A final interesting research direction is system recovery after a 51% attack on a coin chain.
8. Acknowledgements
This work was supported by Ministry of Education (MOE) Singapore’s Tier 2 Grant Award No. MOE-T2EP20120-0003.
Appendix A Proofs of Security Theorems
@@
Before we prove the main theorems, we need the following technical lemmas. We need the first to prove that if all CDPs are overcollateralized, then so is system as a whole.
Lemma A.1.
@@
Proof.
Note that $\frac{\sum_{i=1}^{n}a_{i}}{\sum_{i=1}^{n}b_{i}}=\sum_{i=1}^{n}\frac{a_{i}}{b_{i}}\frac{b_{i}}{\sum_{j=1}^{n}b_{j}}$. Since $\sum_{i=1}^{n}\frac{b_{i}}{\sum_{j=1}^{n}b_{j}}=1$, $\frac{\sum_{i=1}^{n}a_{i}}{\sum_{i=1}^{n}b_{i}}$ is a convex combination of $a_{i}/b_{i},\ldots,a_{n}/b_{n}$. As such, it cannot be smaller than $\min_{i=1,\ldots,n}a_{i}/b_{i}$, which is greater than $c$ by the statement of the lemma. ∎
@@
Proof of Theorem 4.1.
If $p_{y_{j}\,t}c_{j\,t}/y_{j\,t}>\gamma$ holds for all $j\in\mathcal{J}$, then by Lemma A.1 it must hold that
(3) $\frac{C^{+}_{0}}{D^{+}_{0}}>\gamma.$
@@
(4) $D^{-}_{0}<\zeta^{\prime}/(1-\zeta^{\prime})D^{+}_{0}.$
@@
where the first inequality holds because of (4) and the second because of (3). ∎
Proof of Theorem 4.2.
@@
(5) $D^{-}_{0}+h^{\prime}<\zeta^{\prime}/(1-\zeta^{\prime})D^{+}_{0}$
@@
where the first inequality holds because of (5) and the second because $C^{-}_{0}>0$ and because of (3). As such, we again find that the debt ceilings ensure that CroCoDai remains fully backed if it is sufficiently overcollateralized. ∎
Proof of Theorem 4.3.
In the following, we first assume that $O$ is odd and that some subset $O^{\prime}\subset O$ is controlled by malicious parties who aim to convince a vault in CroCoDai that the price is at most than some $p^{*}>p$. In this case, the median equals the $\frac{1}{2}(M+1)+|O^{\prime}|$ th order statistic from a sample of $M-|O|$ normally distributed random variables with mean $p$ and standard deviation $\sigma_{o}^{2}$. This order statistic is greater than the maximum of the $|M|-|O^{\prime}|$ random variables, which we denote by $Y$. Let $\sigma_{*}=\min_{o\in O}\sigma_{o}$. $Y$’s cumulative distribution function equals
$\begin{split}\mathbb{P}(Y\geq p^{*})&=\mathbb{P}\left(\bigwedge_{o\in O\backslash M^{\prime}}\Delta_{o}>p^{*}-p\right)=\prod_{o\notin O^{\prime}}R\left(\frac{p^{*}-p}{\sigma_{o}}\right)\\ &\leq R^{|M|-|O|}\left(\frac{p^{*}-p}{\sigma_{*}}\right)\leq\frac{e^{-(|M|-|O^{\prime}|)(p^{*}-p)^{2}/2}}{(p^{*}-p)\sigma\sqrt{2\pi}},\end{split}$
where the first inequality holds because $R$ is a monotonically decreasing function, and the second inequality follows from Lemma A.2. As such, the probability $\mathbb{P}(Y\geq p^{*})$ vanishes exponentially as $p^{*}-p$ increases. The proof for $p^{*}<p$ is very similar and uses the symmetry of the normal distribution. ∎
Appendix B DSS Parametrization
The parameters of the various different vault types in the DSS can be found here. A sample of parameter values as of 25 April 2023 (Ethereum block 17121982) can be found in Table 5. For each of the supported token type $j$, we indicate the token name, total value of the collateral of CDPs of this type $\sum_{j\in\mathcal{J}_{t}}p_{j\,t}c_{j\,t}$, total debt of the collateral of CDPs of this type $\sum_{j\in\mathcal{J}_{t}}D_{j\,t}$, collateralization ratio, and debt ceiling $\zeta_{j}$. Although the exact values vary throughout the day, this table is meant to illustrate the typical composition of the collateral type in the DSS.